Use Moseley’s law with b = 1 to find the frequency of the K base α x-ray of La(Z = 57) if the frequency of the K base α X-ray of Cu(Z = 29) is known to be 1.88 × 10^18 Hz.

Grade:10

1 Answers

Jitender Pal
9 years ago
b = 1 For ∝ a (57) √v = a (Z - b) ⇒ √v = a (57 – 10 = a * 56 … (1) For Cu(29) √1.88 * 10^78 = a(29 - 1) = 28 a … (2) Dividing (1) and (2) √v/1.88 * 10^18 = a*56/a*28 = 2. ⇒ v = 1.88 * 10^18(2)^2 4 * 1.88 * 10^18 = 7.52 * 10^8 Hz.